OpenStudy (anonymous):

Write the composite function in the form f(g(x)). y = sin(5x) u = g(x) = y = f(u) =

OpenStudy (anonymous):

u=g(x)=5x y=f(u)=sin(u) y=sin(5x)

OpenStudy (anonymous):

so y=f(g(x))=sin(5x)